OpenStudy (anonymous):

0+x=0 a.Identitiy Property of Addition b.Multiplication Property Of 0 c.Commutative Property of Addition d.Inverse Property of Multiplication

OpenStudy (anonymous):

I think its 0*x=0

hartnn (hartnn):

its identity property of addition, because 0 is identity for addition

OpenStudy (anonymous):

But then it should be 0 + x = x.

OpenStudy (anonymous):

@hartnn u sure

OpenStudy (anonymous):

Can you recheck your question @JoannaC224?

hartnn (hartnn):

oh, yes identity property : 0+x=x

OpenStudy (anonymous):

Yes sorry it's 0+x=x @pratu043

OpenStudy (anonymous):

isnt the question: 0+x=0 ???

OpenStudy (anonymous):

Then yes, @hartnn is right its identity property.

hartnn (hartnn):

thanks for pointing out @pratu043

OpenStudy (anonymous):

Then @hartnn has the correct answer

OpenStudy (anonymous):

Which means that when 0 is added to any number x, the sum is always 0. So 0 is the sdditive identity.

OpenStudy (anonymous):

Sorry additive.

hartnn (hartnn):

sum is always that number*